The sum of A and B is 35. Ten years ago A was double B's age. How old are they now?
Answer by JulietG(1812)

A + B = 35, which means that A = 35-B
(A-10) = 2(B-10)
Distribute
A-10 = 2B-20
Substitute the value of A from the first equation
(35-B)-10 = 2B-20
25-B = 2B - 20
Add 20 to each side
45 - B = 2B
Add B to each side
45 = 3B
Divide each side by 3
15 = B
If B is 15, then A must be 20.
.
Let's check...
Ten years ago, A would have been 10 and B would have been 5 -- 10 is double 5.
Success!